A DC offset needs to be applied with a voltage summer to add 1.65 V to the amplified ECG signal. The 1.65 V needs to come from 3.3 V using a voltage divider. The voltage divider output needs to be buffered to the voltage summer. Design the circuit to for the adding of DC offset.
Answer #1

Answer:

This circuit can be made by cascading an op-amp based unity-gain voltage follower circuit followed by a non-inverting voltage adder. The voltage follower acts like a buffer, thus preventing the loading of voltage divider by the voltage summer stage. This design is shown in the image below:

A note on non-inverting adder:

The following circuit adds two voltages (V1 and V2) without inverting their polarities.

If the two input resistance are equal, then the relation between V1, V2 and Vout is given by

Vout = (1+Rf/R1)(V1+V2)/2

So if we keep Rf = R1 (as we have done in case of our design) then the above expression reduces to

Vout = V1+V2

Hope this helps.